If the length of vector a is 5 units and that of b is 4 units and the angles are theta1=38 and theta2=39 degrees, what is the length of the resultant vector?
Answer #1

angle between vector A and B,

x = 39 - 38 = 1

A = 5

B = 4

resultant ve tor

R^2 = A^2 + B^2 + 2 A B cos x

R^2 = 5^2 + 4^2 + 2* 5*4* cos 1

R = 8.99997 units.

=======

do comment in case any doubt, will reply for sure.. Goodluck